# Conceptualization of this study, Methodology, Software 2]organization=Academy of Mathematics and Systems Science, Chinese Academy of Sciences, city=Beijing, postcode=100190, country=China

### v1

@@ Abstract
Recent advances on time series forecasting mainly focus on improving the forecasting models themselves. However, managing the length of the input data can also significantly enhance prediction performance. In this paper, we introduce a novel approach called Optimal Starting Point Time Series Forecast (OSP-TSP) to capture the intrinsic characteristics of time series data. By adjusting the sequence length via leveraging the XGBoost and LightGBM models, the proposed approach can determine optimal starting point (OSP) of the time series and thus enhance the prediction performances. The performances of the OSP-TSP approach are then evaluated across various frequencies on the M4 dataset and other real-world datasets. Empirical results indicate that predictions based on the OSP-TSP approach consistently outperform those using the complete dataset. Moreover, recognizing the necessity of sufficient data to effectively train models for OSP identification, we further propose targeted solutions to address the issue of data insufficiency.
@@ 1 Introduction
Time series forecasting plays a crucial role in practical applications across various fields including services, tourism, finance, meteorology and many others. Before the rise of machine learning — particularly deep learning — traditional forecasting methods predominantly relied on statistical models like Exponential Smoothing (ETS) Hyndman et al. (2008) and ARIMA Box et al. (2015). These models, which use past observations to construct linear functions for predicting future trends, have been widely employed in forecasting tasks for decades. In recent years, deep learning techniques, such as Recurrent Neural Networks (RNNs) Lipton (2015) and ConvTrans Li et al. (2019), have gained popularity in time series forecasting. These approaches excel at capturing complex nonlinear patterns in the data, leading to more accurate predictions.
@@
Previous literature on time series modeling has mainly concentrated on model selection and optimization, such as enhancing forecast accuracy through the integration of forecasting, ensemble learning, and artificial neural networks. These techniques rely on learning from the entire dataset to make predictions. Typically, providing a predictive model with all available data allows it to form a comprehensive understanding of historical patterns, thereby enhancing the model’s predictive capabilities. With recent advances in data storage technologies, we can now easily store and retrieve vast amounts of time series data for model training. However, when structural changes occur in the data, using the full dataset may lead to predictions that place too much weight on long-term trends and potentially overlooking recent shifts that have emerged after these changes.
@@
For example, the flow of tourists between China and Japan had been steadily increasing since 1979, with a significant surge after 2010. However, the outbreak of COVID-19 in 2020 had a profound global impact, and the subsequent quarantine measures imposed by various countries caused a sharp decline in travel. As a result, the number of Chinese tourists visiting Japan in 2020 and 2021 plummeted to levels reminiscent of the pre-2010 era. Currently, tourism is in a phase of recovery, gradually returning to the levels observed before 2019. If we use total tourist numbers data for future predictions, it might overly emphasize the consistent growth observed before 2019, failing to account for the tourism industry’s recovery from the pandemic. This could lead to significant inaccuracies in the forecast. While certain time series models, such as ETS and LSTM, give more weight to recent data to minimize the noise from long-term trends, it may be more appropriate in this case to exclude data prior to the COVID-19 outbreak to improve forecast accuracy. However, simply truncating data may not always result in the most optimal forecast. In fact, we should expect that there exists a specific point in the time series such that if we begin the forecast from that point, the forecast error can be minimized. We refer to this as the optimal starting point (OSP) of the time series. In the previous example, we subjectively chose the starting point of a new time series based on the onset of the COVID-19 pandemic. However, for other time series data, we may not have clear prior knowledge, as we do with tourism data, to pinpoint the location of structural changes that would inform data truncation. Moreover, there are various ways to truncate a time series, making it difficult to identify this point subjectively. Therefore, our goal is to develop a machine learning model that can help us automatically determine the OSP of a time series. This is the primary focus of this article.
@@
It is worth noting that different time series are likely to vary in length, which brings challenge for model training. When feeding a set of time series into a model, the length inconsistency can complicate the training process. To address this issue, we consider an approach similar to change point detection, where we extract consistent features from each time series before putting them into the model for training. These features, such as length, frequency, and other relevant metrics, provide important information into the characteristics of the time series and play a crucial role in improving forecasting accuracy. In existing research, time series features are often directly input into prediction models to enhance accuracy. For instance, models like LightGBM Mamonov et al. (2022) have been successfully applied to time series forecasting. Additionally, time series features are also used in meta-learning for combinatorial forecasting, where they help assign weights to different base models. For example, the FFORMA (Feature-based Forecast Model Averaging) method, introduced by Montero-Manso et al. (2021), uses 42 statistical features (implemented via the R package tsfeatures) to estimate optimal weights for nine traditional models, trained using XGBoost. In this study, we will use LightGBM and XGBoost models learn the characteristics of the time series and determine the location of the optimal starting point.
@@ 3.2 Feature Extraction
Table 1: Number of M4 series per data frequency and domain
@@ 3.3 Training the Optimal Starting Point Model
  • XGBoost XGBoost is an efficient implementation of the Gradient Boosting Decision Tree (GBDT) and offers improvements over the original GBDT Chen and Guestrin (2016). As a forward addition model, its core strategy is to ensemble multiple weak learners into a strong one using boosting. The output of each decision tree is the difference between the target value and the prediction result of all preceding trees. The final result is obtained by summing up all these outputs:
@@
    where $y_{i}^{(t)}$ is the prediction result for sample i after the $t$th, $y_{i}^{(t-1)}$ is the prediction results for the perecious $t-1$ trees and $f_{t}(x_{i})$ is the $t$th DT.
    XGBoost is an efficient implementation of GBDT in that it successfully alleviates the overfitting problem by introducing a regularization function $\Omega(f)$:
@@
  • LightGBM
    LightGBM is a framework that implements the Gradient Boosting Decision Tree (GBDT) algorithm. It supports efficient parallel training and offers advantages such as faster training speed, lower memory consumption, improved accuracy, and distributed support (Ke et al. (2017)). The primary motivation behind the development of LightGBM was to address the challenges faced by GBDT in handling massive data. It aims to accelerate the training of GBDT models without compromising accuracy, thereby enabling GBDT to be used more effectively and swiftly in industrial applications.
    Initially, LightGBM introduces a leaf-wise growth strategy with a depth limit. This strategy avoids indiscriminate treatment of leaves at the same level, which can reduce more errors and achieve better accuracy with the same number of splits, while avoiding unnecessary overhead.
    Furthermore, the Gradient-based One-Side Sampling (GOSS) algorithm adopts a sample reduction approach. It excludes most samples with small gradients and calculates the information gain using only the remaining samples. This approach provides a balanced algorithm in terms of reducing the amount of data and ensuring accuracy.
    High-dimensional data tend to be sparse, and this sparsity has inspired the design of a lossless method to reduce the dimensionality of features. As such, the Exclusive Feature Bundling (EFB) algorithm posits that the number of features can be reduced if certain features are fused and bound together.
@@ 3.5.2 MAPE
4 Empirical analysis
In this section, we present several empirical analysis to demonstrate the finite sample performances of the proposed OSP-TSP algorithm. Code for reproducing the results is available at https://github.com/feng-li/forecasting-with-optimal-starting-point.
@@ 4.1 The M4 dataset
We have chosen the M4 dataset to validate the effectiveness of our proposed Optimal Starting Point (OSP) model construction process. The M4 dataset, compiled by the National Technical University of Athens (NTUA), comprises a total of 100,000 time series collected from several publicly accessible websites, ensuring a diverse and rich set of data sources. Furthermore, the M4 dataset spans multiple industry sectors, including services, tourism, imports and exports, demographics, education, government, bonds, stocks, insurance, loans, real estate, transportation, and natural resources and environment Spiliotis et al. (2020). This wide coverage lends strong representation and applicability to the field of time series forecasting.The specific data types in various fields are shown in Table 2
In addition, the M4 dataset boasts a large capacity, encompassing six data frequencies (annual, quarterly, monthly, weekly, daily, and hourly) and six application areas (micro, macro, financial, demographic, and others). It’s worth noting that, on average, the sequences in the M4 dataset are longer than those in the M3 dataset. This provides more opportunities for complex methods that require substantial amounts of data for effective training Kang et al. (2017). The forecast steps are determined based on the nature of the decisions that the data frequency is most likely to support within a company or organization. For instance, annual data is typically used to support long-term strategic decisions spanning one to five years into the future. Conversely, high-frequency data is typically used to support short-term operational decisions, ranging from a few hours to a few weeks. Makridakis et al. (2020).
@@
Initially, we select three types of seasonal data (annual, quarterly, and monthly) that have the largest volume for empirical research. For each type of data, 70% of the data is selected as the training set to train the Optimal Starting Point (OSP) prediction model. In this context, the ETS and thetaf models are used as the basic prediction model $M$. The sequence segmentation parameters $m$ and $n$ are set to 5 and 4, respectively, and MASE is used as the minimum target error to train the optimal starting section. The trained OSP model is then used to predict the optimal starting points for the remaining data. These predicted optimal starting points are used to reconstruct the new data and make predictions using ETS and thetaf. We then output the average error of the improved prediction of the reconstructed sequence using the optimal starting point in the test data and compare it with the average error of the prediction using the full data.
Table 3 displays the prediction errors MAPE and MASE on the test set corresponding to the eight optimal starting point prediction methods (the combination of the three conditions mentioned in Chapter 1). MASE, as our training target, directly reflects the prediction improvement, while MAPE error indicates whether other non-target errors can be concurrently improved in this process. From Table 3, we observe that the MASE error predicted by almost all the data modified with the optimal starting point is smaller than that predicted by the complete sequence, indicating that the error improvement effect has been achieved. Regarding the improved results of MAPE error, we note that Improved prediction accuracy is also achieved in most cases.
Table 3: OSP improved predictions for ETS and thetaf models on M4 data.
@@
If we were to construct eight different improvement methods to improve the forecast for the same data, it would render the program overly complex. Therefore, we consider opting for only a subset of these improvement methods. The ‘mean’ column in the aforementioned table represents the result of training with the minimum average error as the target. It can be observed that training with the average minimum error as the objective consistently yields better results than training with the absolute minimum. This can be attributed to two factors. Firstly, the instability that arises from intercepting the change of prediction error at the starting point of the time series. Secondly, our final prediction is constructed by taking the four equally divided points of the optimal starting interval of the prediction, and then averaging the values after prediction. This approach is also similar to the idea of minimum average error, which requires the overall prediction results to be better within an interval. Consequently, we can attempt to use only the lowest average error as a prediction.
Acknowledging that the subinterval division may be coarse, we divided each sequence into an average of 10 subintervals ($m=10$), also setting $n=4$, and subsequently output the forecast improvement of our model across the three frequencies. The error results are presented in Table 4. It can be observed that the forecast results, when using the optimal prediction starting point under this model setting, are still significantly superior to those obtained using the complete dataset. Concurrently, we compared the forecast results of the model with those set to be divided into 5 sub-intervals. We found that the model with the finer molecular interval did not yield a significant improvement effect. This suggests that employing more sub-intervals may not necessarily lead to better results, and that the use of fewer sub-intervals can also yield good prediction effects and reduce our workload in the model training phase.
Table 4: Prediction results of OSP model at $m=10$.
@@
In order to demonstrate the universality of our improvement method across different frequency data, we continue to evaluate the improved error forecast results based on ETS and thetaf for weekly, daily, and hourly data in the M4 dataset. This time, we only use the method of marking the interval with the minimum average error to train the prediction results after training the model. The results are presented in Table 5. It can be observed that for the improved target MASE, the improved results of all models exhibit higher prediction accuracy than those obtained using the complete dataset. Concurrently, MAPE also shows a certain degree of improvement in some models.
Simultaneously, we have distilled and presented the five most pivotal variables within the framework of various seasonal models, as detailed in Table 6. Notably, regardless of the seasonal model in question, Curvature and Linearity occupy paramount positions, forming the cornerstone features for predicting optimal starting points. For datasets exhibiting pronounced seasonality, $seas\_acf1$ (seasonal autocorrelation at lag 1) and $seasonal\_strength$ consistently rank high across quarterly, monthly, and even hourly models, emphatically demonstrating the indispensable nature of seasonal factors across these diverse timescales.
Furthermore, we observe that the autocorrelation of time series, as another crucial characteristic, manifests itself in differing forms contingent upon the frequency of the data. In low-frequency annual data, the autocorrelation coefficients of the raw series, particularly $x\_acf1$ (autocorrelation at lag 1) and $x\_acf10$ (sum of squares of the first ten autocorrelation coefficients), exhibit heightened significance. However, as the data frequency escalates to weekly, daily, and even hourly levels, the autocorrelation features of differenced series, such as $diff1\_acf1$ (autocorrelation at lag 1 after first-order differencing) and $diff2\_acf1$ (autocorrelation at lag 1 after second-order differencing), become increasingly pivotal. In summary, for more granular time series data, the autocorrelation features of series after multiple differencing emerge as the decisive factors in identifying and determining optimal starting points.
Table 5: Error results for weekly, daily, and hourly data in M4 data.
@@
Table 6: The importance of the top 5 features under different seasonal models.
@@
4.2 Real world data
Given that the M4 dataset is an artificially constructed dataset, and although we know it originates from various industries in reality, it’s challenging for us to identify its exact source. Therefore, we resort to using data from more transparent sources for empirical study. We selected some data for testing, and the experiment was conducted on five real-world datasets - GDP, Construction Industry, Exchange Rate, Confidence Index, and Import Value, detailed in the appendix. We also employ metrics such as Mean Absolute Scaling Error (MASE) and Mean Absolute Percentage Error (MAPE) to evaluate the performance of our proposed approach. Each dataset is partitioned into a training set and a test set, adhering to a 70/30 split, where 70% of the data is allocated for training and the remaining 30% is reserved for testing. We train the model on the training set, and then use the trained model on the remaining dataset to predict the optimal starting point. Subsequently, we construct a new time series to improve the prediction results. The prediction results on the dataset are output as follows. It can be observed that improvements on these datasets using the optimal starting point of prediction of our model consistently yield better prediction results. Meanwhile, to reflect the adaptability of our model’s prediction, it can theoretically be applied to any basic prediction method. Therefore, we supplement the prediction results of ARIMA and nnetar-based prediction models, and we can observe that our use of the optimal starting point can also achieve better prediction results in these two models.
Table 7: The prediction results of the real dataset via OSP under four basic prediction methods.

### v2

@@ Abstract
Recent advances on time series forecasting mainly focus on improving the forecasting models themselves. However, when the time series data suffer from potential structural breaks or concept drifts, the forecasting performance might be significantly reduced. In this paper, we introduce a novel approach called Optimal Starting Point Time Series Forecast (OSP-TSP) for optimal forecasting, which can be combined with existing time series forecasting models. By adjusting the sequence length via leveraging the XGBoost and LightGBM models, the proposed approach can determine the optimal starting point (OSP) of the time series and then enhance the prediction performances of the base forecasting models. To illustrate the effectiveness of the proposed approach, comprehensive empirical analysis have been conducted on the M4 dataset and other real world datasets. Empirical results indicate that predictions based on the OSP-TSP approach consistently outperform those using the complete time series dataset. Moreover, comparison results reveals that combining our approach with existing forecasting models can achieve better prediction accuracy, which also reflect the advantages of the proposed approach.
@@ 1 Introduction
Time series forecasting plays a crucial role in practical applications across various fields including services, tourism, finance, meteorology and many others. Before the rise of machine learning — particularly deep learning — traditional forecasting methods predominantly relied on statistical models like Exponential Smoothing (ETS) (Hyndman et al., 2008) and ARIMA (Box et al., 2015). These models, which use past observations to construct linear functions for predicting future trends, have been widely employed in forecasting tasks for decades. In recent years, deep learning techniques, such as Recurrent Neural Networks (RNNs) (Lipton, 2015) and ConvTrans (Li et al., 2019), have gained popularity in time series forecasting. These approaches excel at capturing complex nonlinear patterns in the data, leading to more accurate predictions.
@@
Previous literature on time series modeling has mainly concentrated on model selection and optimization, such as enhancing forecast accuracy through the integration of forecasting, ensemble learning, and artificial neural networks. These techniques rely on learning from the entire dataset to make predictions. Typically, providing a predictive model with all available data allows it to form a comprehensive understanding of historical patterns, thereby enhancing the model’s predictive capabilities. With recent advances in data storage technologies, we can now easily store and retrieve vast amounts of time series data for model training. However, in many practical applications, the distribution and trends of data are likely to change over time. There may be sudden shifts at specific time points or gradual, continuous changes over time. These phenomena are referred to as structural breaks or concept drift (Gama et al., 2014; Žliobaitė et al., 2016) in machine learning. Therefore, if these changes in data are not taken into account, the predictive accuracy of the model might be significantly reduced.
@@
To illustrate the idea, consider a real example in tourism. The flow of tourists between China and Japan had been steadily increasing since 1979, with a significant surge after 2010. However, the outbreak of COVID-19 in 2020 had a profound global impact, and the subsequent quarantine measures imposed by various countries caused a sharp decline in travel. As a result, the number of Chinese tourists visiting Japan in 2020 and 2021 plummeted to levels reminiscent of the pre-2010 era. Currently, tourism is in a phase of recovery, gradually returning to the levels observed before 2019. If we use total tourist numbers data for future predictions, it might overly emphasize the consistent growth observed before 2019, failing to account for the tourism industry’s recovery from the pandemic. This could lead to significant inaccuracies in the forecast. While certain time series models, such as ETS and LSTM, give more weight to recent data to minimize the noise from long-term trends, it may be more appropriate in this case to exclude data prior to the COVID-19 outbreak to improve forecast accuracy.
However, simply truncating the data might not necessarily yield to optimal predictive results. In fact, we should expect that there exists a specific point in the time series such that if we begin the forecast from that point, the forecast error can be minimized. We refer to this as the optimal starting point (OSP) of the time series. In the previous example, we subjectively chose the starting point of a new time series based on the onset of the COVID-19 pandemic. However, for other time series data, we may not have clear prior knowledge, to pinpoint the location of structural changes that would inform data truncation. Moreover, there are various ways to truncate a time series, making it difficult to identify this point subjectively. Therefore, our goal is to develop a general framework that can help us automatically determine the OSP of a time series, especially for time series with structural breaks or concept drifts. This is the primary focus of this article.
@@
In this paper, we propose a novel approach called Optimal Starting Point Time Series Forecast (OSP-TSP) for optimal time series forecasting. The proposed OSP-TSP method contains two main steps. In the first step, the proposed method capture the intrinsic characteristics of time series data. Then the proposed method can determine an OSP via leveraging some forecasting models. In this paper, we use the XGBoost and LightGBM (Mamonov et al., 2022) models. In the second step, existing forecasting models can be applied to achieve the final prediction results. It is worth noting that different time series are likely to vary in length, which brings challenge for model training. When feeding a set of time series into a model, the length inconsistency can complicate the training process. To address this issue, we consider extracting consistent features from each time series before putting them into the model for training. These features, such as length, frequency, and other relevant metrics, provide important information into the characteristics of the time series and play a crucial role in improving forecasting accuracy. The performances of the OSP-TSP approach are then evaluated across various frequencies on the M4 dataset and other real-world datasets. Empirical results indicate that predictions based on the OSP-TSP approach consistently outperform those using the complete dataset. Moreover, comparison results reveals that combining our approach with some existing forecasting models can achieve better prediction accuracy, which also reflect the advantages of the proposed method.
@@ 2.1 Optimal starting point determination
The proposed method is also closely related to the field of concept drift as we mentioned previously. Researchers have proposed several solutions in the past literature to address concept drift in machine learning and time series forecasting. According to Liu et al. (2023), previous efforts can be mainly categorized into two directions. The first research direction is to detect structural breaks or concept drifts, and then process them to enhance the predictive accuracy of time series models. An intuitive idea is to combine the change point detection techniques with different time series forecasting models, see for example (Hadad et al., 2017; Wan et al., 2024; Oh and Han, 2000). Another approach is to construct a concept drift detector or filter. Once the prediction metrics reach an alert level, the detector will prompt the model to retrain using the buffered contextual data; see for example (Widmer and Kubat, 1996; Gonçalves Jr and De Barros, 2013). The second research direction focuses on handling concept drifts in an implicit manner. These methods typically aim to establish robust forecasting models that can withstand the effects of concept drift. For instance, Kolter and Maloof (2007) proposed an ensemble learning framework, which can handle concept drift by dynamically adjusting and weighting the base learners in the candidates pool. Liu et al. (2023) considered establishing a forecasting model based on two subsequences, which dynamically adjusts the weights of sequence samples.
@@ 3.2 Feature Extraction
Table 1: Features used for time series data
@@ 3.3 Training the Optimal Starting Point Model
  • XGBoost is an efficient implementation of the Gradient Boosting Decision Tree (GBDT) and offers improvements over the original GBDT (Chen and Guestrin, 2016). As a forward addition model, its core strategy is to ensemble multiple weak learners into a strong one using boosting. The output of each decision tree is the difference between the target value and the prediction result of all preceding trees. The final result is obtained by summing up all these outputs:
@@
    where $y_{i}^{(t)}$ is the prediction result for sample i after the $t$-th, $y_{i}^{(t-1)}$ is the prediction results for the perecious $t-1$ trees and $f_{t}(x_{i})$ is the $t$-th DT. XGBoost successfully alleviates the overfitting problem by introducing a regularization function $\Omega(f)$:
@@
  • LightGBM is a framework that implements the Gradient Boosting Decision Tree (GBDT) algorithm. It supports efficient parallel training and offers advantages such as faster training speed, lower memory consumption, improved accuracy, and distributed support (Ke et al., 2017). The primary motivation behind the development of LightGBM was to address the challenges faced by GBDT in handling massive data. It aims to accelerate the training of GBDT models without compromising accuracy, thereby enabling GBDT to be used more effectively and swiftly in industrial applications. Initially, LightGBM introduces a leaf-wise growth strategy with a depth limit. This strategy avoids indiscriminate treatment of leaves at the same level, which can reduce more errors and achieve better accuracy with the same number of splits, while avoiding unnecessary overhead. Furthermore, the Gradient-based One-Side Sampling (GOSS) algorithm adopts a sample reduction approach. It excludes most samples with small gradients and calculates the information gain using only the remaining samples. This approach provides a balanced algorithm in terms of reducing the amount of data and ensuring accuracy. High-dimensional data tend to be sparse, and this sparsity has inspired the design of a lossless method to reduce the dimensionality of features. As such, the Exclusive Feature Bundling (EFB) algorithm posits that the number of features can be reduced if certain features are fused and bound together.
@@ 3.5.2 MAPE
4 Numerical Studies
In this section, we present several numerical studies to demonstrate the finite sample performances of the proposed OSP-TSP algorithm. Codes for reproducing the results are available at https://github.com/feng-li/forecasting-with-optimal-starting-point.
@@ 4.1 The M4 dataset
We have chosen the M4 dataset to validate the effectiveness of our proposed Optimal Starting Point (OSP) model construction process. The M4 dataset, compiled by the National Technical University of Athens (NTUA), comprises a total of 100,000 time series collected from several publicly accessible websites, ensuring a diverse and rich set of data sources. Furthermore, the M4 dataset spans multiple industry sectors, including services, tourism, imports and exports, demographics, education, and many others (Spiliotis et al., 2020). This wide coverage lends strong representation and applicability to the field of time series forecasting. The specific data types in various fields are shown in Table 2. Moreover, the M4 dataset is distinguished by its substantial size and comprehensive coverage, including six different data frequencies—annual, quarterly, monthly, weekly, daily, and hourly—as well as six application domains: microeconomic, macroeconomic, financial, demographic, and others. Notably, the time series in the M4 dataset are, on average, longer than those in the M3 dataset, offering more extensive data for complex methodologies for effective training (Kang et al., 2017).
@@
To start with, we select three types of seasonal data (i.e. yearly, quarterly, and monthly) that have the largest volume for empirical research. For each type of data, 70% of the data is selected as the training set to train the OSP prediction model. Specifically, the ETS and thetaf models are chosen as the basic forecasting model $M$. We consider four different improvement methods under two different class for each forecasting model. This leads to a total of sixteen improvement methods. The sequence segmentation parameters $m$ and $n$ are set to be 5 and 4, respectively. We use the MASE metric as the target error to be minimized for training the OSP model. The trained OSP model is then used to predict the optimal starting points for the remaining data. These predicted optimal starting points are used to reconstruct the new data and make predictions using ETS and thetaf models. The averaged error (i.e. MAPE and MASE) of the prediction results based on OPS-TSP method on test data are reported in Table 3, as well as those results that using the total time series data. Here, the MASE results, as our training target, directly reflect the prediction improvement, while the MAPE results indicate whether other non-target errors can be improved simultaneously. The best results are presented in boldface.
From Table 3, we find that the MASE error results obtained using our OSP-TSP method are smaller in almost all cases as compared to those prediction results obtained by using the total time series. This indicates that our method indeed yields better prediction results. Similar results can also be found in most cases in terms of MAPE results. Moreover, it would render the program overly complex if we construct eight different improvement methods to improve the prediction results. Therefore, we consider opting for only a subset of these improvement methods. The ‘mean’ column in Table 3 represents the results of training with the minimum average error as the target. We find that training with the average minimum error as the objective consistently yields better results than training with the absolute minimum. This can be attributed to two factors. First, the instability that arises from intercepting the change of prediction error at the starting point of the time series. Second, our final prediction is constructed by taking the four equally divided points of the optimal starting interval of the prediction, and then averaging the values after prediction. This approach is also similar to the idea of minimum average error, which requires the overall prediction results to be better within an interval. Consequently, we can attempt to use only the lowest average error as a prediction.
Table 3: Prediction error results for yearly, quarterly, and monthly data with $(m,n)=(5,4)$.
@@
Next, noticing that the sub-interval division may be coarse, We consider a more refined division that dividing each sequence into an average of $m=10$ sub-intervals, while keeping $n=4$. The other settings are similar to the previous numerical studies. The prediction error results are presented in Table 4. One can observe that the prediction results obained by our OSP-TSP method are still superior to those obtained using the total time series. In the meanwhile, compared with the prediction results in Table 3, we find that the model with finer molecular intervals does not yield a significant improvement effect. This suggests that employing more sub-intervals might not necessarily lead to better results. Consequently, the use of fewer sub-intervals can also yield good prediction effects and reduce our workload in the model training phase.
Table 4: Prediction error results for yearly, quarterly, and monthly data in M4 data with $(m,n)=(10,4)$
@@
In order to demonstrate the universality of our improvement method across different frequency data, we further extend our numerical studies for weekly, daily, and hourly data in the M4 dataset. For simplicity, we only use the method of marking the interval with the minimum average error to train the prediction results after training the model. The results are presented in Table 5. We have observed similar prediction results, which indecates the universal effectiveness of the proposed method.
[…]
Table 5: Prediction error results for weekly, daily, and hourly data in M4 data with $(m,n)=(5,4)$
@@
Table 6: The importance of the top 5 features under different seasonal data.
@@
4.2 Other real world data
In this subsection, we consider five more real world datasets to demonstrate the finite sample performances of the proposed OSP-TSP approach. Specifically, they are GDP, Construction Industry, Exchange Rate, Confidence Index, and Import Value. As we mentioned previously, the proposed OSP-TSP approach is a general framework, which can be integrated with a variety of time series forecasting models. In this numerical experiment, we consider the ARIMA model and the neural network time series forecasting model nnetar as our base forecasting model. We still employ metrics such as Mean Absolute Scaling Error (MASE) and Mean Absolute Percentage Error (MAPE) to evaluate the prediction results of our proposed approach. Each dataset is partitioned into two parts, where 70% of the data is allocated for training and the remaining 30% is reserved for testing. We first use the training set to train the OSP model, and then use the trained model on the remaining dataset to predict the optimal starting point. Subsequently, we construct a new time series to improve the prediction results. The prediction results on the five real world datasets are reported in Table 7. One can find that the OSP-TSP approach consistently yield better prediction results than those using the total time series data under the same base models in most cases. It is worth noting that we find some models do not perform well on the exchange rate dataset. In fact, under the efficient market hypothesis, one would expect a random walk to perform best on these series, which only uses one data point. Therefore, the proposed OSP-TSP approach does not lead to consistent improvements.
Table 7: Prediction error results for five real world datasets with four different base models. The best results are in boldface.